OpenStudy (anonymous):

4y^2/3y^2-6y+3 x 3y-3/2y solve. i got as far as 12y^3-12y^2 / 6y^3-12y^2+6y and now am stuck

OpenStudy (anonymous):

I have no idea what equation to work with.

OpenStudy (anonymous):

factor the second numerator and the first denominator first\[=\frac{4y^2}{3(y-1)^2}*\frac{3(y-1)}{2y}\]the y-1 expression will cancel; the 3's will cancel; the 4 and the 2 will reduce; and the y^2 and the ywill reduce\[=\frac{2y}{y-1}\]

OpenStudy (anonymous):

did i skip too many steps

OpenStudy (anonymous):

nope :) thank you!

OpenStudy (anonymous):

cool! :})